### v1

@@ Abstract
keywords:
[…]
@@ 1 Introduction
In recent years, the rapid development of deep learning techniques has injected new momentum into emerging directions in econometrics and opened up possibilities for addressing long-standing challenges. Current research has largely focused on deep learning methods themselves, tackling issues such as high-dimensional data, complex models, and unstructured data. For instance, Mancini (2021), Maliar et al. (2021), and Feng et al. (2024) use deep learning to handle high-dimensional prediction, solve dynamic economic models, and construct deep factor models, respectively. Applications to unstructured data, as in Dell (2025), demonstrate the capability of deep learning to extract meaningful economic signals from large-scale text or image. Hybrid modeling approaches have also been developed, integrating deep neural networks into traditional models. Studies such as Liu et al. (2023), Stok et al. (2024), and Chen et al. (2024) exemplify the fusion of deep learning with traditional econometric structures to enhance both interpretability and model performance. From a theoretical perspective, the integration of econometric methods with deep learning offers new potential for methodological advancement, particularly in embedding deep learning more seamlessly into traditional econometric frameworks.
@@
Here, we focus on the Granger causality introduced by Granger (1969), which provides a practical and mathematical framework for analyzing dynamic relationships between time series. Since then, Granger causality has been widely applied in various fields from natural to social science (Becker and Woessmann, 2013; Seth et al., 2015; Abatis et al., 2024). Predictability is a key feature of Granger causality analysis, making it highly relevant to economists, policymakers, and investors. Extensive research has exploited Granger causality to uncover interrelationships among economic and financial variables. Notable applications include the detection of financial risk Hong et al. (2009); Balboa et al. (2015); Corsi et al. (2018), as well as broader economic analyzes Puente-Ajovín and Sanso-Navarro (2015); Ramos-Herrera and Sosvilla-Rivero (2025).
@@
Real-world data often exhibit complex underlying structures (Tong, 1990; Fan and Yao, 2003; Granger, 2004), making the accurate characterization of nonlinear causality a key challenge. Therefore, it is preferable that Granger causality methods are not constrained by rigid model specifications. As a result, nonparametric approaches to the Granger causality test have received increasing attention. Hidalgo (2000) introduced a nonparametric causality test in the frequency domain for weakly stationary linear processes, particularly for long-range-dependent observations. Su and White (2008) proposed a nonparametric test for conditional independence based on the weighted Hellinger distance between conditional densities, which can be directly applied to testing Granger causality. Nishiyama et al. (2011) proposed a nonparametric test based on moment conditions, which allows for nonlinear Granger causality. Song and Taamouti (2018) developed model-free measures for Granger causality in mean and derived the asymptotic distribution of the associated nonparametric estimator, making it applicable for testing causality. Nonparametric tests are highly flexible, particularly in detecting nonlinear forms of causality. However, in practice it can be difficult to determine the appropriate lag length for real data. In economics and finance, the effects of variables often take time to manifest—for instance, the impact of economic policies or financial regulations may only become apparent after a considerable delay. As the lag order in Granger causality tests increases, the performance of nonparametric methods tends to deteriorate due to the curse of dimensionality. They may fail to control the type I error or suffer from low power against relevant alternatives. Consequently, their practical applicability may be limited, and the development of methods that are less sensitive to dimensionality remains an important research challenge.
@@
Due to its strong capacity to approximate complex nonlinear relationships, particularly its exceptional performance in high-dimensional settings, deep learning has attracted increasing attention from researchers. In recent years, theoretical advances have begun to emerge regarding the statistical inference of deep learning models. Semiparametric inference for deep neural networks is crucial to make our proposal feasible. Farrell et al. (2021) established the convergence rate of deep ReLU networks for bounded i.i.d. data. Zhou et al. (2023) extended similar results to data that satisfy the $\beta$ mixing conditions, particularly for the mixture density network (MDN) introduced by Bishop (1994). Brown (2024) further proposed the convergence rate of deep ReLU networks for unbounded dependent data. Our proposal relies on a multilayer perceptron (MLP) to estimate the conditional mean, and an MDN to effectively estimate conditional density distribution. Both components play an important role in the construction of the test statistic. Our test requires convergence rate guarantees for the MLP and the MDN based on semiparametric inference theories.
@@
One should be aware that directly applying deep neural network estimators in the classic nonparametric test will lead to inevitable issues, such as excessive bias and inability to control the type I error. The root cause of these problems lies in the lack of established convergence rate guarantees for deep neural networks, makes it impossible to ensure that the underlying stochastic process converges at an adequate rate. This, in turn, hinders the application of limit theorems for empirical processes in constructing the test statistic. To overcome this difficulty, we adopt a doubly robust structure, originally introduced by Robins et al. (1994). The doubly robust framework has been extensively employed in causal inference and missing data problems, as it ensures consistency provided that at least one of the involved models is correctly specified. See Bang and Robins (2005) and Vermeulen and Vansteelandt (2015) for example. Chernozhukov et al. (2018) were among the first to introduce the doubly robust principle into the context of machine learning. More recently, Zhou et al. (2023) developed a novel testing framework by incorporating the doubly robust principle into the study of Markov properties. Unlike in the missing data setting, when both models are well specified, the doubly robust approach in testing can further accelerate the convergence rate. For Granger causality, previous studies have explored its use with deep neural networks, also known as neural Granger causality (e.g., Tank et al., 2021; Khanna and Tan, 2019). These methods identify the Granger causality by imposing group sparsity penalties on the weights of the first hidden layer. Although neural Granger causality methods provide a convenient approach to uncovering causal relationships among a large number of time series, they may be sensitive to hyperparameter choices and training procedures and lack formal statistical guarantees.
@@
Our proposal contributes to the literature in several meaningful ways. The most immediate advantage lies in practical applications. In real data analysis, selecting the appropriate lag order for Granger causality tests is often challenging, as researchers may have little guidance. The proposed DRGCT mitigates this issue by leveraging the strong performance of deep neural network estimators when handling multiple lag orders. As a result, one can choose a sufficiently large lag length that encompasses all potential causal effects, thereby gaining a more comprehensive understanding of the underlying causal relationships. In addtion, we introduce deep learning techniques into the framework of the Granger causality test in a novel way. To apply deep neural network estimators within a classical statistical testing framework, we construct a doubly robust hypothesis structure to test Granger causality. We obtain a sharp upper bound on the approximation error for the MLP estimator under the $\beta$ mixing condition for both bounded and unbounded data based on Farrell et al. (2021) and Brown (2024). By combining the convergence rate of the MDN estimator as in Zhou et al. (2023), we demonstrate that the proposed test successfully controls type I error and achieves power approaching one. Notably, we establish theoretical validity for testing with unbounded data through deep learning–based estimators. In this sense, our work extends the applicability of hypothesis testing by incorporating deep learning tools. Furthermore, our test statistic achieves a parametric convergence rate while its components are estimated semiparametrically. The difference in results between the doubly robust statistic and the single-estimator statistic reflects its effect on controlling the type I error. For local power analysis, tests based on empirical processes can achieve a parametric rate of $n^{-1/2}$, which is much faster than the rates obtained from smoothing-based nonparametric tests. It is worth noting that our proposed test does not rely on sample splitting or cross-fitting, as it effectively controls the type I error without these techniques. This offers the advantage of using the entire dataset for model training, thus improving the efficiency of estimation. Finally, thanks to its doubly robust structure, the proposed test can employ a multiplier bootstrap method to obtain the critical value. Unlike the standard wild bootstrap, this procedure allows the critical value to be determined using the already estimated quantities, thereby avoiding redundant calculations. This feature is particularly important for deep learning–based methods, as it substantially reduces computational cost and running time.
@@
The rest of the article is organized as follows. Section 2 provides the motivation for the doubly robust test and introduces the corresponding hypothesis and the test statistic. Section 3 presents the main theoretical results of the proposed test. Specifically, Section 3.1 outlines the necessary lemmas that establish the convergence rates of the relevant deep neural network estimators, providing the theoretical foundation for our proposed test and reflecting the advantage of doubly robust construction when the convergence rates of individual estimators are insufficient. Section 3.2 establishes the asymptotic distribution of the test statistic under the null hypothesis, while Section 3.3 discusses the consistency of the test and its asymptotic local power under local alternatives. In Section 3.4, we present a bootstrap procedure for determining the critical value and implementing the testing procedure. Section 4 reports the simulation results, including a comparison with the nonparametric test in Nishiyama et al. (2011), and Section 5 illustrates the methodology using real data sets. Section 6 is the conclusion. All assumptions and technical proofs are provided in Appendix.
@@ 2 Hypotheses and test statistic
We depart from the settings in the classical nonlinear Granger causality test by allowing for large lag orders $p$ and $q$. Traditional nonparametric methods for estimating $m(\boldsymbol{Y}_{t-1})$ may fail due to "the curse of dimensionality". Therefore, we apply deep learning techniques, which are particularly well suited to estimate ${m}(\boldsymbol{Y}_{t-1})$ in this situation. However, from Chernozhukov et al. (2018), naively applying a deep neural network to estimate ${m}(\boldsymbol{Y}_{t-1})$ would cause a heavy bias, making $\widehat{S}_{n}^{0}$ in (5) not have a tractable limiting distribution under the null hypothesis.
@@
The reorganized samples $\{Y_{t},\boldsymbol{Y}_{t-1}\}_{t=q+1}^{n}$ are used to train the MLP and obtain conditional expectation estimators $\left\{\widehat{m}(\mathbf{Y}_{t-1})\right\}_{t=q+1}^{n}$, where $\boldsymbol{Y}_{t-1}$ and $Y_{t}$ are the input covariates and the response variable, respectively. For estimating $\phi(\nu|\boldsymbol{Y}_{t-1})$, we begin with inputting $\{\boldsymbol{X}_{t-1},\boldsymbol{Y}_{t-1}\}_{t=q+1}^{n}$ to train the estimated conditional density function $\widehat{f}_{\boldsymbol{X}_{t-1}|\boldsymbol{Y}_{t-1}}(x|y)$ through the MDN framework, where $\boldsymbol{Y}_{t-1}$ and $\boldsymbol{X}_{t-1}$ are the input and response, respectively. For each $t\in\{q+1,q+2,\cdots,n\}$, given $\boldsymbol{Y}_{t-1}$, we randomly generate $M$ samples $\left\{\boldsymbol{X}_{j}^{*}\right\}_{j=1}^{M}$ by $\widehat{f}_{\boldsymbol{X}_{t-1}|\boldsymbol{Y}_{t-1}}(x|y)$. Let $\widehat{\phi}\left(\nu|\boldsymbol{Y}_{t-1}\right)=\frac{1}{M}\sum_{j=1}^{M}\exp\left(i\nu^{\top}\boldsymbol{X}_{j}^{*}\right)$. Then, we obtain $\left\{\widehat{\phi}\left(\nu|\boldsymbol{Y}_{t-1}\right)\right\}_{t=q+1}^{n}$ for some $\nu\in\mathcal{W}_{2}$. For details on the MDN framework, see Bishop (1994) or Zhou et al. (2023). The estimation procedures are summarized as Step 1 and Step 2 in Algorithm 1.
@@
Input: Data $\{X_{t},Y_{t}\}_{t=1}^{n}$, the corresponding numbers of lags $p$ and $q$, the number of pairs $L$, the number of samples from the generators $M$, the number of bootstrap replications $B$, and the significance level $\alpha$.
@@
(2b) Use $\boldsymbol{Y}_{t-1}$ as input covariates and $\boldsymbol{X}_{t-1}$ as the response variable to train the MDN model and obtain the estimated conditional density $\widehat{f}_{\boldsymbol{X}_{t-1}|\boldsymbol{Y}_{t-1}}(x|y)$.
@@
(2e) Calculate the conditional characteristic function estimators $\widehat{\phi}\left(\nu_{l}|\boldsymbol{Y}_{t-1}\right)=\frac{1}{M}\sum_{m=1}^{M}\exp\left(i\nu_{l}^{\top}\boldsymbol{X}_{m}^{*}\right)$ for $l=1,\cdots,L$ and $t=q+1,\cdots,n$.
@@ 3 Theory
In this section, we first discuss the convergence rates of deep neural network based estimators $m(\boldsymbol{Y}_{t-1})$ and $\phi(\nu|\boldsymbol{Y}_{t-1})$, which serve as the theoretical basis for the proposed test. Then, we present the asymptotic properties of the empirical processes (8) and the test statistic $\mathrm{KS}_{n}$ (10) under the null and alternatives in Section 3.2 and 3.3, respectively. Finally, we obtain the critical value of the test through a multiplier bootstrap procedure and verify its validity.
@@ 3.1 Convergence rates of deep neural networks
Suppose Assumptions 1, 3 and 4 hold. Suppose the MLP is specified with width $H_{n}\asymp n^{\left(\frac{q}{\beta_{0}+q}\right)\left(1/2-\kappa_{B}\right)}\log^{2}(n)$ and depth $L_{n}\asymp\log(n)$. Let $\epsilon_{n}=n^{-\left(\frac{\beta_{0}}{\beta_{0}+q}\right)\left(1/2-\kappa_{B}\right)}\log^{6}(n)$. Then, for the MLP estimator $\widehat{m}(\boldsymbol{Y}_{t-1})$, there exist constants $C,C_{1},C_{2}>0$ independent of $n$, for all $n$ sufficiently large, such that
$P\left(\mathbb{E}\left|\widehat{m}(\boldsymbol{Y}_{t-1})-m(\boldsymbol{Y}_{t-1})\right|^{2}\leq C^{2}\epsilon_{n}^{2}\right)\geq 1-e^{-n\left(\frac{\beta_{0}}{\beta_{0}+q}\right)^{\left(1/2-\kappa_{B}\right)}}-\frac{2C_{1}n^{1-C_{2}\log(n)}}{\log(n)}-o(1/n),$
$P\left(\mathbb{E}_{n}\left|\widehat{m}(\boldsymbol{Y}_{t-1})-m(\boldsymbol{Y}_{t-1})\right|^{2}\leq C^{2}\epsilon_{n}^{2}\right)\geq 1-4e^{-n\left(\frac{\beta_{0}}{\beta_{0}+q}\right)^{\left(1/2-\kappa_{B}\right)}}-\frac{12C_{1}n^{1-C_{2}\log(n)}}{\log(n)}-o(1/n).$
@@
Suppose Assumption 1 and 5 hold. Then, there exists a certain MDN function class satisfying Assumption 6, such that the resulting MDN estimator $\widehat{f}_{\boldsymbol{X}_{t-1}|\boldsymbol{Y}_{t-1}}(x|y)$ satisfies that, with probability at least $1-O\left(n^{-1}\right)$,
@@
for some constants $C,\omega_{1},\omega_{2}>0$.
@@
Suppose Assumption 1 and 7 (ii)-(iv) hold. Then, there exists a constant $\kappa>1/2$, such that

### v2

@@ Abstract
Keywords: Granger causality; Deep learning; Empirical processes; Multiplier bootstrap; Hypothesis test
@@ 1 Introduction
Due to it’s ability to process and analyze large complex datasets, the rapid development of deep learning has injected new momentum into economic and financial studies. The applications of deep learning range from classical demand estimation (Chernozhukov et al., 2019) and solving dynamic economic models (Maliar et al., 2021) to risk management and volatility prediction (Sadhwani et al., 2021; Liu et al., 2023; Stok et al., 2024), asset pricing (Nagel, 2021; Gu et al., 2021; Chen et al., 2024) and portfolio allocation (Caner and Daniele, 2022). Deep learning has proven effective in tackling high-dimensional nonlinear modeling problems, which is also a key concern for econometricians. Pioneers have actively explored the value of deep learning in econometric theory (Farrell et al., 2021; Chernozhukov et al., 2022; Hyun and Wang, 2025), especially the potential to address long-standing theoretical challenges.
@@
In this study, we focus on the Granger causality introduced by Granger (1969), which is a mathematical framework with significant practical guidance to analyze dynamic relationships between time series. White and Lu (2010) provided rigorous structural characterizations of Granger causality to avoid misuse of Granger’s pioneering concept. Predictability is a key feature in the Granger causality analysis and thus draws enormous attention from financial market participants and economic policy makers. Applications include the detection of financial risk (Hong et al., 2009; Balboa et al., 2015; Corsi et al., 2018), as well as extensive research on the discovery of interrelationships between economic variables (Becker and Woessmann, 2013; Puente-Ajovín and Sanso-Navarro, 2015; Ramos-Herrera and Sosvilla-Rivero, 2025). White and Pettenuzzo (2014) presented a specific framework for economic policy analysis. Granger causality has also been widely applied in other various fields, from natural science (Shojaie and Michailidis, 2010; Seth et al., 2015; Abatis et al., 2024) to social science (Soysa and Oneal, 1999; Seetanah, 2011; Gunter and Önder, 2016).
@@
It is preferable that Granger causality methods are not constrained by a rigid model specification, since real-world data often exhibit complex nonlinear relationships (Tong, 1990; Fan and Yao, 2003; Granger, 2004). Therefore, nonparametric approaches to the Granger causality test have received increasing attention. Hidalgo (2000) introduced a nonparametric causality test in the frequency domain for weakly stationary linear processes, particularly for long-range-dependent observations. Su and White (2008) proposed a nonparametric test for conditional independence based on the weighted Hellinger distance between conditional densities, which can be directly applied to testing Granger causality. Nishiyama et al. (2011) proposed a nonparametric test based on moment conditions, which allows nonlinear Granger causality. Song and Taamouti (2018) developed model-free measures for Granger causality in mean and derived the asymptotic distribution of the associated nonparametric estimator, making it applicable for testing causality.
Although nonparametric tests are highly flexible in detecting nonlinear Granger causality, the lag orders have to be chosen very small, which is a fundamental issue inherent in present methods. However, in most economic and financial application scenarios, the effects of a set of variables on others often take a while to work out. For example, the impact of economic policies or financial regulations may only become apparent after a considerable delay (Friedman, 1961; Jovanovski and Muric, 2011; Aruoba and Drechsel, 2024). As the lag orders in Granger causality tests increase, traditional nonparametric methods become ineffective even fail due to the curse of dimensionality. Considering the strong capacity to approximate complex nonlinear relationships, particularly exceptional performance in high-dimensional settings, neural network based learning for Granger causality is invented (e.g., Tank et al., 2021; Nauta et al., 2019; Khanna and Tan, 2020). Learning for Granger causality opens up a new perspective for Granger causality study; however, these methods may not be suitable for econometric analysis, since it lacks support for statistical inference.
@@
In recent years, theoretical advances have begun to emerge regarding the statistical inference of deep learning models. Farrell et al. (2021) established the convergence rate of deep ReLU networks for bounded i.i.d. data. Zhou et al. (2023) extended similar results to data that satisfy the $\beta$ mixing conditions, particularly for the mixture density network (MDN) introduced by Bishop (1994). Brown (2024) further proposed the convergence rate of deep ReLU networks for unbounded dependent data. Our proposal relies on a multilayer perceptron (MLP) to estimate the conditional mean, and an MDN to effectively estimate conditional density distribution. Both components play an important role in the construction of the test statistic. Our test requires convergence rate guaranties for the MLP and the MDN based on semiparametric inference theories.
@@
One should be aware that directly applying deep neural network estimators in the classic nonparametric test will lead to inevitable issues, such as excessive bias and the inability to control type I error. The root cause of these problems lies in the lack of established convergence rate guaranties for deep neural networks, making it impossible to ensure that the underlying stochastic process converges at an adequate rate. This, in turn, hinders the application of limit theorems to empirical processes in constructing the test statistic. To overcome this difficulty, we adopt a doubly robust structure, originally introduced by Robins et al. (1994). The doubly robust framework has been extensively employed in causal inference and missing data problems, as it ensures consistency, provided that at least one of the involved models is correctly specified. See Bang and Robins (2005) and Vermeulen and Vansteelandt (2015) for example. Chernozhukov et al. (2018) were among the first to introduce the doubly robust principle into the context of machine learning. More recently, Zhou et al. (2023) developed a novel testing framework by incorporating the doubly robust principle into the study of Markov properties. Unlike in the missing data setting, when both models are well specified, the doubly robust approach in testing can further accelerate the convergence rate.
@@
Our proposal contributes to the literature in several meaningful ways. The most immediate advantage lies in practical applications. In real data analysis, selecting the appropriate lag order for Granger causality tests is often challenging, as researchers may have little guidance. The proposed DRGCT mitigates this issue by leveraging the strong performance of deep neural network estimators when handling multiple lag orders. As a result, one can choose a sufficiently long lag that encompasses all potential causal effects, thereby gaining a more comprehensive understanding of the underlying causal relationships. In addition, we introduce deep learning techniques into the framework of the Granger causality test in a novel way. To apply deep neural network estimators within a classical statistical testing framework, we construct a doubly robust hypothesis structure to test Granger causality. We obtain a sharp upper bound on the approximation error for the MLP estimator under the $\beta$ mixing condition for both bounded and unbounded data based on Farrell et al. (2021) and Brown (2024). By combining the convergence rate of the MDN estimator as in Zhou et al. (2023), we demonstrate that the proposed test successfully controls type I error and achieves a power approaching one. In particular, we establish theoretical validity for testing with unbounded data using deep learning–based estimators. In this sense, our work extends the applicability of hypothesis testing by incorporating deep learning tools. Furthermore, our test statistic achieves a parametric convergence rate, while its components are estimated semiparametrically. The difference in results between the doubly robust statistic and the single-estimator statistic reflects its effect on controlling the type I error. For local power analysis, tests based on empirical processes can achieve a parametric rate of $n^{-1/2}$, which is much faster than the rates obtained from smoothing-based nonparametric tests. It is worth noting that our proposed test does not rely on sample splitting or cross-fitting, as it effectively controls the type I error without these techniques. This offers the advantage of using the entire dataset for model training, thus improving the efficiency of estimation. Finally, thanks to its doubly robust structure, the proposed test can employ a multiplier bootstrap method to obtain the critical value. Unlike the standard wild bootstrap, this procedure allows the critical value to be determined using the already estimated quantities, thereby avoiding redundant calculations. This feature is particularly important for deep learning–based methods, as it substantially reduces computational cost and running time.
@@
The remainder of the article is organized as follows. Section 2 provides the motivation for the doubly robust test and introduces the corresponding hypothesis and the test statistic. Section 3 presents the main theoretical results of the proposed test. Specifically, Section 3.1 outlines the necessary lemmas that establish the convergence rates of the relevant deep neural network estimators, providing the theoretical foundation for our proposed test and reflecting the advantage of doubly robust construction when the convergence rates of individual estimators are insufficient. Section 3.2 establishes the asymptotic distribution of the test statistic under the null hypothesis, while Section 3.3 discusses the consistency of the test and its asymptotic local power under local alternatives. In Section 3.4, we present a bootstrap procedure for determining the critical value and implementing the testing procedure. Section 4 reports the simulation results, including a comparison with the nonparametric test in Nishiyama et al. (2011), and Section 5 illustrates the methodology using real data sets. Section 6 is the conclusion. All assumptions and technical proofs are provided in Appendix.
@@ 2 Hypotheses and test statistic
We depart from the settings in the classical nonlinear Granger causality test by allowing for large lag orders $p$ and $q$. Traditional nonparametric methods for estimating $m(\boldsymbol{Y}_{t-1})$ may fail due to ”the curse of dimensionality”. Therefore, we apply deep learning techniques, which are particularly well suited to estimate ${m}(\boldsymbol{Y}_{t-1})$ in this situation. However, from Chernozhukov et al. (2018), A naive application of a deep neural network to estimate ${m}(\boldsymbol{Y}_{t-1})$ will cause a substantial bias in the test statistic based on (5) under the null hypothesis.
@@
The reorganized samples $\{Y_{t},\boldsymbol{Y}_{t-1}\}_{t=q+1}^{n}$ are used to train the MLP and obtain conditional expectation estimators $\left\{\widehat{m}(\mathbf{Y}_{t-1})\right\}_{t=q+1}^{n}$, where $\boldsymbol{Y}_{t-1}$ and $Y_{t}$ are the input covariates and the response variable, respectively. For estimating $\phi(\nu|\boldsymbol{Y}_{t-1})$, we begin with inputting $\{\boldsymbol{X}_{t-1},\boldsymbol{Y}_{t-1}\}_{t=q+1}^{n}$ to train the estimated conditional density function $\widehat{f}_{\boldsymbol{X}_{t-1}|\boldsymbol{Y}_{t-1}}(x|y)$ by the MDN with $G$ mixture components, where $\boldsymbol{Y}_{t-1}$ and $\boldsymbol{X}_{t-1}$ are the input and response, respectively. For each $t\in\{q+1,q+2,\cdots,n\}$, given $\boldsymbol{Y}_{t-1}$, we randomly generate $M$ samples $\left\{\boldsymbol{X}_{j}^{*}\right\}_{j=1}^{M}$ by $\widehat{f}_{\boldsymbol{X}_{t-1}|\boldsymbol{Y}_{t-1}}(x|y)$. Let $\widehat{\phi}\left(\nu|\boldsymbol{Y}_{t-1}\right)=\frac{1}{M}\sum_{j=1}^{M}\exp\left(i\nu^{\top}\boldsymbol{X}_{j}^{*}\right)$. Then, we obtain $\left\{\widehat{\phi}\left(\nu|\boldsymbol{Y}_{t-1}\right)\right\}_{t=q+1}^{n}$ for some $\nu\in\mathcal{W}_{2}$. For details on the MDN framework, see Bishop (1994) or Zhou et al. (2023). The estimation procedures are summarized as Step 1 and Step 2 in Algorithm 1.
@@
Input: Data $\{X_{t},Y_{t}\}_{t=1}^{n}$, the corresponding numbers of lags $p$ and $q$, the number of mixture components of the MDN $G$, the number of pairs $L$, the number of samples from the generators $M$, the number of bootstrap replications $B$, and the significance level $\alpha$.
@@
(2b) Use $\boldsymbol{Y}_{t-1}$ as input covariates and $\boldsymbol{X}_{t-1}$ as the response variable to train the MDN model with $G$ mixture components and obtain the estimated conditional density $\widehat{f}_{\boldsymbol{X}_{t-1}|\boldsymbol{Y}_{t-1}}(x|y)$.
@@
(2e) Calculate the conditional characteristic function estimators $\widehat{\phi}\left(\nu_{l}|\boldsymbol{Y}_{t-1}\right)=\frac{1}{M}\sum_{j=1}^{M}\exp\left(i\nu_{l}^{\top}\boldsymbol{X}_{j}^{*}\right)$ for $l=1,\cdots,L$ and $t=q+1,\cdots,n$.
@@ 3 Theory
In this section, we first discuss the convergence rates of the deep neural network based estimators $\widehat{m}(\boldsymbol{Y}_{t-1})$ and $\widehat{\phi}(\nu|\boldsymbol{Y}_{t-1})$, which serve as the theoretical basis for the proposed test. Then, we present the asymptotic properties of the empirical processes (8) and the test statistic $\mathrm{KS}_{n}$ (10) under the null and alternatives in Section 3.2 and 3.3, respectively. Finally, we obtain the critical value of the test through a multiplier bootstrap procedure and verify its validity.
@@ 3.1 Convergence rates of deep neural networks
Suppose Assumptions 1, 3 and 4 hold. Suppose the MLP is specified with width $H_{n}\asymp n^{\left(\frac{q}{\beta_{0}+q}\right)\left(1/2-\kappa_{B}\right)}\log^{2}(n)$ and depth $L_{n}\asymp\log(n)$. Let $K_{n}=n^{-\left(\frac{\beta_{0}}{\beta_{0}+q}\right)\left(1/2-\kappa_{B}\right)}\log^{6}(n)$. Then, for the MLP estimator $\widehat{m}(\boldsymbol{Y}_{t-1})$, there exist constants $C,C_{1},C_{2}>0$ independent of $n$, for all $n$ sufficiently large, such that
$P\left(\mathbb{E}\left|\widehat{m}(\boldsymbol{Y}_{t-1})-m(\boldsymbol{Y}_{t-1})\right|^{2}\leq C^{2}K_{n}^{2}\right)\geq 1-e^{-n\left(\frac{\beta_{0}}{\beta_{0}+q}\right)^{\left(1/2-\kappa_{B}\right)}}-\frac{2C_{1}n^{1-C_{2}\log(n)}}{\log(n)}-o(1/n),$
$P\left(\mathbb{E}_{n}\left|\widehat{m}(\boldsymbol{Y}_{t-1})-m(\boldsymbol{Y}_{t-1})\right|^{2}\leq C^{2}K_{n}^{2}\right)\geq 1-4e^{-n\left(\frac{\beta_{0}}{\beta_{0}+q}\right)^{\left(1/2-\kappa_{B}\right)}}-\frac{12C_{1}n^{1-C_{2}\log(n)}}{\log(n)}-o(1/n).$
@@
Suppose Assumptions 1 and 5 hold. Then, there exists a certain MDN function class satisfying Assumption 6, such that the resulting MDN estimator $\widehat{f}_{\boldsymbol{X}_{t-1}|\boldsymbol{Y}_{t-1}}(x|y)$ satisfies that
@@
for some constants $C,\omega_{1},\omega_{2}>0$, with probability at least $1-O\left(n^{-1}\right)$.
@@
Suppose Assumptions 1 and 7 (ii)-(iv) hold. Then, there exists a constant $\kappa>1/2$, such that